.. 4c1/5c2 = 2/5 ..Anonymous wrote:I wrote down this problem, but forgot to write down the answer.
5 apples and 1 of them is spoiled. If two apples are pulled out, what is the probability that one of them will be spoiled?
I solved it like this:
The prob(no spoils) = 4/5*3/4 = 12/20 = 3/5
1 - prob(no spoils) = at least one spoil = 1 - 3/5 = 2/5
Does anyone else get 2/5 as the correct answer?
Thanks!
GMATLiveTeachStarts Nov 22
LIVE ONLINE CLASSES
Get Ready for GMAT Test Day Faster with Live Online Classes
with Josh Braslow, 20+ years in test prep
- Nov 22 · Josh Braslow
- Sun · 9:30 AM to 12:30 PM ET
- Included
- 40 hours of live online classes + 6 months of TTP OnDemand
- Attend the first class for free
- Every class is recorded, so you never fall behind













